NOAA-15, HIRS, Filter Wheel Housing Heater Turned On, 19 April 2004

5:23 PM
Subject: NOAA-15, HIRS, Filter Wheel Housing Heater Turned On, 19 April 2004
- Issued: 19 April 2004; 2115
To: undisclosed-recipients:;

HIRS increased noise due to filter wheel motor current surge; heater
turned on

Date/Time(UTC) Message Issued:
19 April 2004, 2115 UTC

Satellite(s) Involved:

Instrument(s) Involved:

Product(s) Involved/Affected:
TBD, noise in data

Date/Time(UTC) of Initial Implementation:
19 April 2004, Time 1250 UTC

Details/Specifics of Change:

SOCC is reporting the following:
N15 HIRS filter wheel motor current, jumped suddenly by 20 mA at 01:35 UTC
today (day 110). Over the next several hours the current rose more
gradually to as high as 302 mA. At 02:05 UTC the period monitor began to
show excursions beyond its nominal range and at 03:13 UTC the filter wheel
began to lose sync intermittently. During later passes the motor current
and temperature, as well as the housing temperature began to flag. To begin
the correction process, the filter heater was turned on during the 12:50 UTC
pass. It is planned to leave it on until the motor current comes down and
the synch word and period monitor become more stable. This usually takes a
few days. Users will see increased noise in the data. SOCC and IPD will
continue monitoring.

To minimize the impact on the level 1B data, normal updates of the 24 hour
calibration file have been suspended to prevent contamination. Also, lunar
intrusion detection will be turned off to prevent incorrect flagging of the

Contact Person(s) Name/Email/Phone Number for Questions: ext. 196 instrument scientist ext. 129 with ops questions

Web site(s) that contain other relevant information (where applicable): - History of level 1B
notices - All user notices - SOCC polar satellites status - SOCC morning reports